# MedGRPO: Multi-Task Reinforcement Learning for Heterogeneous Medical Video Understanding

### v1

@@ 3.2 MedGRPO Training Paradigm
We apply GRPO [deepseekmath], a policy gradient method that avoids value function training and naturally handles diverse reward scales through group-relative advantage estimation. For a given prompt , we sample a group of responses from our current policy . We compute the advantage for each response by normalizing the response’s reward within each group:
@@
where is the importance sampling ratio at token position . Following DAPO [dapo], we adopt asymmetric clipping with and to allow larger positive updates while constraining negative ones, and remove the KL penalty term from standard GRPO for improved performance.
@@ 3.3 Cross-Dataset Reward Normalization
Task-Specific Reward Design. We apply GRPO training to four representative tasks spanning different granularities: TAG (segment-level), STG (frame-level), VS (video-level), and RC (segment-level captioning). For temporal grounding (TAG, STG), we use multiplicative composite rewards as penalties rather than additive bonuses, since the model reliably outputs valid structures after SFT: , where is logistic-normalized mIoU and penalizes parsing failures. For captioning (VS, RC), we combine semantic similarity and LLM judge evaluation (§3.4).
@@ 3.4 Medical LLM Judge for Evaluation
For caption generation tasks (VS, DVC, RC), standard embedding-based metrics (SentenceBERT [sentencebert], BERTScore [bertscore]) capture overall paragraph-level semantic similarity but fail to assess fine-grained medical correctness. For example, captions ”The tool grasps tissue in the upper area” vs. ”The grasper dissects the cystic duct in the upper right quadrant” achieve cosine similarity yet differ critically in instrument specificity (generic ”tool” vs precise ”grasper”), action accuracy (”grasps” vs ”dissects”), anatomical precision (vague ”tissue” vs specific ”cystic duct”), and spatial detail (”upper area” vs ”upper right quadrant”). Such distinctions are clinically significant but invisible to general-domain embeddings.
@@
| GPT-4.1 [gpt4v] | 0.018 | 0.250 | 0.087 | 0.014 | 0.096 | 0.005 | 2.438 | 0.101 | 2.490 | 2.080 |
| Gemini-2.5-flash [gemini25] | 0.101 | 0.228 | 0.107 | 0.047 | 0.045 | 0.021 | 2.387 | 0.084 | 2.352 | 1.912 |
| Qwen2.5VL-7B [qwen2.5vl] | 0.105 | 0.151 | 0.010 | 0.020 | 0.006 | 0.068 | 2.512 | 0.075 | 2.452 | 2.090 |
[…]
@@
[…]
LLM Judge Design. We design a GPT-4.1-based judge using comparative similarity scoring: the judge evaluates “How closely does the generated caption match the reference?” rather than rating absolute quality. This comparative framing avoids score inflation [geval] and provides better discrimination between model qualities. The judge evaluates five clinical dimensions with explicit 1–5 rubrics [videochatgpt]: (1) medical terminology precision (clinical terms vs. lay language), (2) instrument & anatomy identification (specific tools and structures), (3) specificity vs. vagueness (precise details vs. generic descriptions), (4) clinical procedure context (workflow and safety awareness), and (5) action & state accuracy (surgical actions and tissue states). For each dimension, scores range from 1 (“completely different”) to 5 (“semantically equivalent”). We compute the mean score and apply logistic normalization: .
@@ 4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL-7B as the base model with frozen vision encoder. During SFT training, video frames are sampled at 0.1-3 FPS with adaptive sampling for longer procedures. During GRPO training, frames are sampled at 1 FPS. Each frame contains 82828 to 482828 pixels. SFT training on the Standard version runs for 3 epochs with learning rate 5e-7 and batch size 8, while scaling law experiments on the Large-Scale version use 1 epoch to demonstrate performance improvements as training data increases from 0 to 461K samples (Figure 4). GRPO training uses 5000 gradient updates with learning rate 5e-7 and batch size 5. We adopt technical improvements from DAPO [dapo] in our GRPO implementation with group size and clipping parameters and . Logistic normalization uses slope and format penalty . All experiments run on 8 H100 GPUs.
@@ 4.2 Main Results
Table 2 presents our main results across all 8 tasks. We compare against off-the-shelf baselines (GPT-4.1, Gemini-2.5-Flash, Qwen2.5VL-7B), our SFT baseline, and our full MedGRPO method (Qwen2.5VL-7B).
Comparison with Off-the-shelf Baselines. Our SFT baseline significantly outperforms all off-the-shelf baselines, demonstrating the importance of domain adaptation. GPT-4.1 achieves only 0.018 on CVS and 0.014 on STG, while Gemini-2.5-Flash reaches 0.101 on CVS and 0.047 on STG. Off-the-shelf Qwen2.5VL-7B performs poorly across grounding tasks (STG: 0.020, TAG@0.3: 0.006). In contrast, our SFT baseline achieves 0.894 on CVS and 0.177 on STG, showing substantial improvements from medical video adaptation, and establishing the importance of a large-scale and diverse benchmark like MedVidBench.
MedGRPO Performance. Our full method (Qwen2.5VL-7B) achieves consistent improvements over SFT across most tasks: CVS (+0.020 to 0.914), STG (+0.025 to 0.202), SA (+0.026 to 0.244), TAG@0.3 (+0.074 to 0.216), TAG@0.5 (+0.065 to 0.156), DVC (+0.132 to 3.797), DVC (+0.045 to 0.210), VS (+0.588 to 4.184), and RC (+0.685 to 3.442). NAP slightly decreases from 0.442 to 0.427, as it was not included in the reward optimization (we use TAG, STG, VS, and RC for computing rewards). Caption generation tasks benefit most from our hybrid LLM judge design, with VS and RC showing substantial improvements. Temporal grounding tasks also show strong gains, with TAG@0.3 and TAG@0.5 improving significantly.
@@ 4.3 Ablation Studies
LLM Judge Design. Rows D-E isolate the contribution of our medical LLM judge for caption generation tasks. Row D (VS+RC with LLM judge) improves VS from 3.596 (SFT baseline) to 3.824 and RC from 2.757 to 3.235, demonstrating that domain-specific evaluation captures medical precision beyond semantic similarity. Row E (VS+RC without LLM judge) achieves only 3.733 on VS and 2.984 on RC, showing the importance of medical-specific evaluation. The full method (Row A) combining all four task rewards achieves the strongest performance: VS 4.184 and RC 3.442, with improvements of 16.4% and 24.8% respectively over SFT baseline.
@@ 5 Conclusion
We present MedVidBench, a large-scale heterogenous benchmark for medical video understanding, and MedGRPO, a multi-dataset RL framework. We find that standard RL collapses on this data due to unbalanced reward scales across datasets. MedGRPO prevents this using two innovations: cross-dataset reward normalization to balance optimization, and a medical LLM judge for domain-specific evaluation. Our method substantially outperforms supervised fine-tuning, establishing that multi-dataset RL in specialized domains requires reward balancing and domain-adapted evaluation.
@@
This supplementary material provides comprehensive details on MedVidBench dataset construction and MedGRPO training methodology. §A describes our data curation pipeline, including prompt design strategies for web-sourced and frame-annotated datasets, QA generation procedures, and human validation study results. §B presents detailed dataset statistics covering task distribution, temporal characteristics, and annotation quality patterns. §C provides complete implementation details for both SFT and GRPO training. §D details our medical LLM judge rubrics across five clinical evaluation dimensions. §E presents additional qualitative results and failure analysis. All materials support reproducibility and provide insights beyond the main paper’s scope constraints.
@@ Appendix A Dataset Curation Pipeline
Our data curation pipeline transforms existing expert annotations from 8 medical video datasets into high-quality instruction-following format using dual multi-modal large language models (GPT-4.1 [gpt4v] and Gemini-2.5-Flash [gemini25]). We employ a multi-perspective approach adapting prompting strategies to dataset characteristics: frame-annoated datasets (CholecT50 [cholect50], EgoSurgery [egosurgery], CholecTrack20 [cholectrack20], CoPESD [copesd]) receive rich contextual prompts incorporating frame-level annotations, while web-sourced datasets (AVOS [avos], NurViD [nurvid]) utilize high-quality audio transcripts (Whisper-X [whisperx]) and video metadata to supplement visual understanding.
@@ A.4 QA Generation and Quality Assurance
After caption generation, we create diverse QA pairs by combining dataset-specific context prefixes with task-specific question templates (3–6 template variants per task), producing instruction-following instances for all 8 tasks spanning video-level, segment-level, and frame-level understanding. To ensure annotation quality, we employ dual-model validation: for caption generation tasks (video summary, dense captioning, region captioning), we independently generate captions using both GPT-4.1 and Gemini-2.5-Flash, compute semantic similarity using sentence-transformers, and filter low-quality pairs with similarity 0.3. This dual-model approach prevents model-specific biases and hallucinations while ensuring consistent high-quality annotations. For evaluating caption quality during both dataset validation and RL training, we design a medical LLM judge (detailed rubrics in §D) that assesses captions across five clinical dimensions through comparative similarity scoring.
@@ A.5 Human Validation Study
To validate our annotation-enriched prompting approach, we conducted a user study with 12 participants who are experts, work in medical data analysis, to compare captions generated using our expert prompts versus a frames-only baseline, both of which are described next. For CoPESD [copesd] dataset, we generated two types of captions: (1) with expert prompt: captions using our full pipeline with overlaid bounding boxes, procedure-specific context, and expert annotations; (2) without expert prompt: captions generated from raw video frames only using a minimal prompt (“Describe what you see in this healthcare procedure video in one sentence”), without procedural context, annotation overlays, timestamps, or domain knowledge. We developed a web interface, shown in Figure 5, to allow participants to rank the caption-pairs. Participants were provided with detailed instructions and examples to select the superior caption based on clinical accuracy and terminological precision. Figure 6 shows the results: participants strongly preferred captions generated with expert prompts (82.0%) over frames-only captions (18.0%), confirming that our annotation-enriched prompting strategy produces superior medical video descriptions compared to naive frame-based generation.
@@ Appendix B MedVidBench: Dataset Statistics
| CholecT50 [cholect50] | Laparoscopic | 50 | ✓ | - | ✓ | - | ✓ | - | ✓ | - | 7.1K |
| CholecTrack20 [cholectrack20] | Laparoscopic | 20 | - | - | - | - | - | ✓ | - | ✓ | 102.7K |
| Cholec80-CVS [cholec80cvs] | Laparoscopic | 80 | - | - | - | ✓ | - | - | - | - | 4.4K |
| CoPESD [copesd] | Laparoscopic | 40 | ✓ | - | ✓ | - | ✓ | ✓ | ✓ | ✓ | 70.3K |
| AVOS [avos] | Open Surgery | 25 | - | - | ✓ | - | ✓ | - | ✓ | - | 62.5K |
| EgoSurgery [egosurgery] | Open Surgery | 21 | - | - | - | - | - | ✓ | - | ✓ | 154.3K |
| JIGSAWS [JIGSAWS] | Robotic Surgery | 103 | - | ✓ | - | - | - | - | - | - | 1.0K |
| NurViD [nurvid] | Nursing | 287 | ✓ | - | ✓ | - | ✓ | - | ✓ | - | 129.5K |
@@
Appendix C Implementation Details
@@
Appendix D Medical LLM Judge: Detailed Rubrics
As described in §3.4 of the main paper, we employ an LLM-as-a-judge approach using GPT-4.1 to evaluate caption quality through comparative assessment across five medical-specific dimensions. Each dimension uses a 1–5 scale measuring how closely the generated caption matches the reference: 5 (very close match, minor phrasing differences), 4 (good match, minor omissions), 3 (partial match, notable omissions), 2 (significant differences, missing important information), 1 (very different, major errors or missing content). We compute the mean score across dimensions and apply logistic normalization (Eq. 3) to obtain the final reward. Below we detail the five evaluation dimensions with scoring rubrics.
Medical Terminology Precision

### v2

@@ 3.2 MedGRPO Training Paradigm
We apply GRPO [45], a policy gradient method that avoids value function training and naturally handles diverse reward scales through group-relative advantage estimation. For a given prompt , we sample a group of responses from our current policy . We compute the advantage for each response by normalizing the response’s reward within each group:
@@
where is the importance sampling ratio at token position . Following DAPO [58], we adopt asymmetric clipping with and to allow larger positive updates while constraining negative ones, and remove the KL penalty term from standard GRPO for improved performance.
@@ 3.3 Cross-Dataset Reward Normalization
Task-Specific Reward Design. We apply GRPO training to four representative tasks spanning different temporal granularities: VS (video-level), TAG (segment-level), RC (segment-level) and STG (frame-level). For grounding tasks (TAG, STG), we use multiplicative composite rewards as penalties rather than additive bonuses, since the model reliably outputs valid structures after SFT: , where is logistic-normalized mIoU and penalizes parsing failures. For captioning tasks (VS, RC), we combine semantic similarity and LLM judge evaluation (§3.4).
@@ 3.4 Medical LLM Judge for Evaluation
For caption generation tasks (VS, DVC, RC), standard embedding-based metrics (SentenceBERT [43], BERTScore [61]) capture overall paragraph-level semantic similarity but fail to assess fine-grained medical correctness. For example, captions “The tool grasps tissue in the upper area” vs. “The grasper dissects the cystic duct in the upper right quadrant” achieve cosine similarity yet differ critically in instrument specificity (generic “tool” vs precise “grasper”), action accuracy (“grasps” vs “dissects”), anatomical precision (vague “tissue” vs specific “cystic duct”), and spatial detail (“upper area” vs “upper right quadrant”). Such distinctions are clinically significant but invisible to general-domain embeddings.
@@
| GPT-4.1 [19] | 0.018 | 0.250 | 0.087 | 0.014 | 0.096 | 0.005 | 2.438 | 0.101 | 2.490 | 2.080 |
| Gemini-2.5-flash [11] | 0.101 | 0.228 | 0.107 | 0.047 | 0.045 | 0.021 | 2.387 | 0.084 | 2.352 | 1.912 |
| VideoChat-R1.5-7B [57] | 0.000 | 0.270 | 0.006 | 0.000 | 0.009 | 0.005 | 1.723 | 0.026 | 3.034 | 3.086 |
| Qwen2.5VL-7B [4] | 0.105 | 0.151 | 0.010 | 0.020 | 0.006 | 0.068 | 2.512 | 0.075 | 2.452 | 2.090 |
| Qwen2.5VL-7B (NVIDIA) [38] | 0.000 | 0.302 | 0.000 | 0.000 | 0.019 | 0.013 | 1.945 | 0.051 | 2.101 | 2.986 |
| Qwen2.5VL-7B (Ours) | 0.894 | 0.442 | 0.218 | 0.177 | 0.142 | 0.091 | 3.665 | 0.165 | 3.596 | 2.757 |
| Qwen2.5VL-7B (Ours) | 0.896 | 0.405 | 0.254 | 0.202 | 0.216 | 0.156 | 3.797 | 0.214 | 4.184 | 3.442 |
@@
| A TAG+STG+VS+RC (full) | 0.896 | 0.405 | 0.254 | 0.202 | 0.216 | 0.156 | 3.797 | 0.214 | 4.184 | 3.442 |
| B TAG+STG+VS+RC w/o normal | 0.020 | 0.267 | 0.234 | 0.010 | 0.004 | 0.003 | 1.061 | 0.002 | 3.805 | 3.469 |
| C TAG+STG w/o VS+RC | 0.914 | 0.394 | 0.257 | 0.193 | 0.202 | 0.142 | 3.718 | 0.225 | 3.776 | 3.425 |
| D VS+RC w/ llm-judge | 0.894 | 0.434 | 0.239 | 0.183 | 0.149 | 0.096 | 3.688 | 0.165 | 3.824 | 3.235 |
| E VS+RC w/o llm-judge | 0.894 | 0.363 | 0.259 | 0.183 | 0.140 | 0.090 | 3.628 | 0.161 | 3.733 | 2.984 |
LLM Judge Design. We design a GPT-4.1-based judge using comparative similarity scoring: the judge evaluates “How closely does the generated caption match the reference?” rather than rating absolute quality. This comparative framing avoids score inflation [31] and provides better discrimination between model qualities. The judge evaluates five clinical dimensions with explicit 1–5 rubrics [34]: (1) medical terminology precision (clinical terms vs. lay language), (2) instrument & anatomy identification (specific tools and structures), (3) specificity vs. vagueness (precise details vs. generic descriptions), (4) clinical procedure context (workflow and safety awareness), and (5) action & state accuracy (surgical actions and tissue states). For each dimension, scores range from 1 (“completely different”) to 5 (“semantically equivalent”). We compute the mean score and apply logistic normalization: .
@@ 4.1 Experimental Setup
Implementation Details. We use Qwen2.5-VL-7B as the primary base model. During SFT training, video frames are sampled at 0.1-3 FPS with adaptive sampling for longer procedures. During GRPO training, frames are sampled at 1 FPS. Each frame contains 82828 to 482828 pixels. SFT training on the Standard version runs for 3 epochs with learning rate 5e-7 and batch size 8, while scaling law experiments on the Large-Scale version use 1 epoch to demonstrate performance improvements as training data increases from 0 to 461K samples (Figure 4). GRPO training uses 5000 gradient updates with learning rate 5e-7 and batch size 5. We adopt technical improvements from DAPO [58] in our GRPO implementation with group size and clipping parameters and . Logistic normalization uses slope and format penalty . All experiments run on 8 H100 GPUs. Additional results with Qwen3VL-4B under both SFT and MedGRPO training are provided in the supplementary.
@@ 4.2 Main Results
Table 2 presents our main results across all 8 tasks. We compare against off-the-shelf baselines (GPT-4.1, Gemini-2.5-Flash, VideoChat-R1.5-7B, Qwen2.5VL-7B, and Qwen2.5VL-7B), our SFT baseline, and our full MedGRPO method (Qwen2.5VL-7B).
Comparison with Off-the-shelf Baselines. Our SFT baseline significantly outperforms all off-the-shelf baselines, demonstrating the importance of domain adaptation. GPT-4.1 achieves only 0.018 on CVS and 0.014 on STG, while Gemini-2.5-Flash reaches 0.101 on CVS and 0.047 on STG. VideoChat-R1.5-7B, a recent RL-trained general-domain VideoLLM, scores 0.000 on CVS and STG, showing that general video RL training does not transfer to medical tasks. Off-the-shelf Qwen2.5VL-7B similarly performs poorly across grounding tasks (STG: 0.020, TAG@0.3: 0.006). Qwen2.5VL-7B, a recent medical VideoLLM fine-tuned on laparoscopic data, also fails to generalize (CVS: 0.000, STG: 0.000), highlighting that single-dataset surgical specialization is insufficient for the diverse tasks in MedVidBench. In contrast, our SFT baseline achieves 0.894 on CVS and 0.177 on STG, establishing the importance of large-scale multi-source training on MedVidBench.
MedGRPO Performance. Our full method (Qwen2.5VL-7B) achieves consistent improvements over SFT across most tasks: CVS (+0.002 to 0.896), STG (+0.025 to 0.202), SA (+0.036 to 0.254), TAG@0.3 (+0.074 to 0.216), TAG@0.5 (+0.065 to 0.156), DVC (+0.132 to 3.797), DVC (+0.049 to 0.214), VS (+0.588 to 4.184), and RC (+0.685 to 3.442). NAP slightly decreases from 0.442 to 0.405, as it was not included in the reward optimization (we use TAG, STG, VS, and RC for computing rewards). Caption generation tasks benefit most from our hybrid LLM judge design, with VS and RC showing substantial improvements. Temporal grounding tasks also show strong gains, with TAG@0.3 and TAG@0.5 improving significantly.
@@ 4.3 Ablation Studies
LLM Judge Design. Rows D-E isolate the contribution of our medical LLM judge for caption generation tasks. Row D (VS+RC with LLM judge) improves VS from 3.596 (SFT baseline) to 3.824 and RC from 2.757 to 3.235, demonstrating that domain-specific evaluation captures medical precision beyond semantic similarity. Row E (VS+RC without LLM judge) achieves only 3.733 on VS and 2.984 on RC, showing the importance of medical-specific evaluation. The full method (Row A) combining all four task rewards achieves the strongest performance: VS 4.184 and RC 3.442, improving 16.4% and 24.8% over SFT baseline.
@@ 5 Conclusion
We present MedVidBench, a large-scale heterogenous benchmark for medical video understanding, and MedGRPO, a multi-dataset RL framework. We find that standard RL collapses on this data due to unbalanced reward scales across datasets. MedGRPO prevents this using two innovations: cross-dataset reward normalization to balance optimization, and a medical LLM judge for domain-specific evaluation. Our method substantially outperforms supervised fine-tuning, establishing that multi-dataset RL requires reward balancing and domain-adapted evaluation.
@@
This supplementary material provides comprehensive details on MedVidBench dataset construction and MedGRPO training methodology. §A describes our data curation pipeline, including prompt design strategies for web-sourced and frame-annotated datasets, QA generation procedures, and human validation study results. §B presents detailed dataset statistics covering task distribution, temporal characteristics, and annotation quality patterns. §C shows additional quantitative results on the baseline model Qwen3-VL. §D provides implementation details for SFT, GRPO training, and skill assessment evaluation. §E details our medical LLM judge rubrics across five clinical evaluation dimensions. §F presents additional qualitative results and failure analysis. All materials support reproducibility and provide insights beyond the main paper’s scope constraints.
@@ Appendix A Dataset Curation Pipeline
Our data curation pipeline transforms existing expert annotations from 8 medical video datasets into high-quality instruction-following format using dual multi-modal large language models (GPT-4.1 [19] and Gemini-2.5-Flash [11]). We employ a multi-perspective approach adapting prompting strategies to dataset characteristics: frame-annoated datasets (CholecT50 [40], EgoSurgery [14], CholecTrack20 [39], CoPESD [53]) receive rich contextual prompts incorporating frame-level annotations, while web-sourced datasets (AVOS [16], NurViD [36]) utilize high-quality audio transcripts (Whisper-X [5]) and video metadata to supplement visual understanding.
@@ A.4 QA Generation and Quality Assurance
After caption generation, we create diverse QA pairs by combining dataset-specific context prefixes with task-specific question templates (3–6 template variants per task), producing instruction-following instances for all 8 tasks spanning video-level, segment-level, and frame-level understanding. To ensure annotation quality, we employ dual-model validation: for caption generation tasks (video summary, dense captioning, region captioning), we independently generate captions using both GPT-4.1 and Gemini-2.5-Flash, compute semantic similarity using sentence-transformers, and filter low-quality pairs with similarity 0.3. This dual-model approach prevents model-specific biases and hallucinations while ensuring consistent high-quality annotations. For evaluating caption quality during both dataset validation and RL training, we design a medical LLM judge (detailed rubrics in §E) that assesses captions across five clinical dimensions through comparative similarity scoring.
@@ A.5 Human Validation Study
To validate our annotation-enriched prompting approach, we conducted a user study with 12 participants who are experts, work in medical data analysis, to compare captions generated using our expert prompts versus a frames-only baseline, both of which are described next. For CoPESD [53] dataset, we generated two types of captions: (1) with expert prompt: captions using our full pipeline with overlaid bounding boxes, procedure-specific context, and expert annotations; (2) without expert prompt: captions generated from raw video frames only using a minimal prompt (“Describe what you see in this healthcare procedure video in one sentence”), without procedural context, annotation overlays, timestamps, or domain knowledge. We developed a web interface, shown in Figure 5, to allow participants to rank the caption-pairs. Participants were provided with detailed instructions and examples to select the superior caption based on clinical accuracy and terminological precision. Figure 6 shows the results: participants strongly preferred captions generated with expert prompts (82.0%) over frames-only captions (18.0%), confirming that our annotation-enriched prompting strategy produces superior medical video descriptions compared to naive frame-based generation.
@@ Appendix B MedVidBench: Dataset Statistics
| CholecT50 [40] | Laparoscopic | 50 | ✓ | - | ✓ | - | ✓ | - | ✓ | - | 7.1K |
| CholecTrack20 [39] | Laparoscopic | 20 | - | - | - | - | - | ✓ | - | ✓ | 102.7K |
| Cholec80-CVS [44] | Laparoscopic | 80 | - | - | - | ✓ | - | - | - | - | 4.4K |
| CoPESD [53] | Laparoscopic | 40 | ✓ | - | ✓ | - | ✓ | ✓ | ✓ | ✓ | 70.3K |
| AVOS [16] | Open Surgery | 25 | - | - | ✓ | - | ✓ | - | ✓ | - | 62.5K |
| EgoSurgery [14] | Open Surgery | 21 | - | - | - | - | - | ✓ | - | ✓ | 154.3K |
| JIGSAWS [1] | Robotic Surgery | 103 | - | ✓ | - | - | - | - | - | - | 1.0K |
| NurViD [36] | Nursing | 287 | ✓ | - | ✓ | - | ✓ | - | ✓ | - | 129.5K |
@@
Appendix C Additional Quantitative Results
Qwen3-VL. Table 5 validates our framework’s generalizability by applying the same SFT and MedGRPO pipeline to Qwen3-VL-4B, a smaller model with improved temporal modeling. Off-the-shelf Qwen3-VL-4B achieves near-zero performance on grounding tasks (STG: 0.000, CVS: 0.000), confirming that architectural advances alone cannot address medical video understanding without domain adaptation. Our SFT training yields strong gains (CVS: 0.895, TAG@0.3: 0.465, TAG@0.5: 0.403), and MedGRPO further improves across all tasks, with notable gains on STG (+0.043), TAG@0.3 (+0.039), and DVC (+0.045). These results demonstrate that MedGRPO generalizes across model architectures and scales, consistently improving upon SFT baselines regardless of the underlying model.
Appendix D Implementation Details
@@
Skill Assessment Evaluation.
We average the 6 OSATS dimension scores and apply thresholds to derive 3 classes (Novice/Intermediate/Expert). We report macro-averaged accuracy (mAcc) and MAE (lower is better): zero-shot Qwen2.5VL-7B achieves MAE=2.440, mAcc=0.000; SFT improves to MAE=1.262, mAcc=0.197; MedGRPO further improves to MAE=1.246, mAcc=0.254, demonstrating the effectiveness of our method.
Appendix E Medical LLM Judge
As described in §3.4 of the main paper, we employ an LLM-as-a-judge approach using GPT-4.1 to evaluate caption quality through comparative assessment across five medical-domain-specific dimensions. Each dimension uses a 1–5 scale measuring how closely the generated caption matches the reference: 5 (very close match, minor phrasing differences), 4 (good match, minor omissions), 3 (partial match, notable omissions), 2 (significant differences, missing important information), 1 (very different, major errors or missing content). §E.1 details how the LLM Judge scores correlates with scores from board certified clinicians on these five dimensions. We describe these five dimensions in details in §E.2 with scoring rubrics.
E.1 Quality Assurance
To establish the validity of our LLM-as-a-judge approach using GPT-4.1 to evaluate caption quality, we conduct a rigorous human study with 10 board-certified clinicians. The clinicians were asked to score 30 samples across the same 5 clinical dimensions used by our LLM Judge (paper lines 404-410). Results show strong correlation: Pearson =, Cohen’s Kappa =, confirming our automated metric effectively proxies human clinical preference. We further show the correlation of the LLM Judge with clinicians across all the five evaluation dimensions (-) in Fig. 8. This highlights that the LLM Judge very closely agrees with clinicians in all of the evaluation dimensions.
E.2 Detailed Rubrics
These five evaluation dimensions used by the LLM Judge and board certified clinicians to score captions is detailed in this section.
Medical Terminology Precision (R1).